Find the square roots of the following numbers by the Prime Factorisation Method.  x. 8100 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

8100 = 2×2×3×3×3×3×5×5×1
⇒ 8100 = (2×2) ×(3×3)×(3×3)×(5×5) ⇒ 8100 = (2×3×3×5)×(2×3×3×5)
⇒ 8100 = 90×90
⇒ 8100 = (90)2
⇒√8100=90


Thanks